# FLIPHAT: Joint Differential Privacy for High Dimensional Sparse Linear Bandits

### v2

@@ Abstract
* signifies shared first authorship.
@@ 1 Introduction
In many of these applications, the personal information used as contexts often has a large dimension, which can be modeled by viewing the contexts as high-dimensional vectors. This is relevant in the modern big data era, where the aforementioned applications usually have many features. However, the established regret lower bound in LCB polynomially depends on the context dimension (Dani et al.,, 2008; Chu et al.,, 2011). Thus, additional structural conditions, such as sparsity, are needed to be enforced to make the problem tractable. Sparsity aligns with the idea that among these features, possibly only a small subset of them are relevant to generate the reward. This setting is called the high-dimensional sparse linear contextual bandit (SLCB). Recently, SLCB has received significant attention (Wang et al.,, 2018; Bastani and Bayati,, 2020; Hao et al.,, 2020; Chakraborty et al.,, 2023) and have been useful in practical applications (Miao and Chao,, 2022; Ren and Zhou,, 2024).
@@
Example 1.
Let us consider a mobile health application that recommends a personalized treatment plan (actions) to each patient (user) based on her personal information, such as age, weight, height, medical history etc. (Tewari and Murphy,, 2017; Killian et al.,, 2023). The recommendations to the users are obtained as feedback through an interactive portal. This can be modeled as a SLCB, such that for every user $u$, context vectors corresponding to each personalized treatment plan $a$ are obtained through some feature representation $x=\phi(u,a)\in\mathbb{R}^{d}$. For example, a feature representation is obtained by inputting the user’s data in a pre-trained neural network. Here, $d$ is often large to capture the complex interactions across treatment plans and user information but there are fewer $s^{*}$ features that finally impact the treatment outcomes. Once the recommendation algorithm provides a treatment plan, it induces a change in the users medical condition. The goal of the recommendation algorithm is to improve the medical condition of the users as much as possible over a number of interactions. Meanwhile, this mobile app’s recommendation algorithm also needs to be updated once in a while in a cloud server after collecting feedback data from a batch of patients. The data includes treatment outcomes and medical contexts of patients, which are sensitive information according to GDPR (Voigt and Von dem Bussche,, 2017) and HIPAA (Annas,, 2003) privacy acts. Hence, each patient would like to obtain a personalized and effective treatment plan while demanding their private information remain protected against any potential adversary.
Related Work: Differential Privacy in Bandits. Problems like the one in Example 1 motivate us to study data-privacy preserving SLCB algorithms. Specifically, we follow Differential Privacy (DP) (Dwork,, 2006) as the data privacy framework. DP aims to ensure that while looking into the output of an algorithm, an adversary is unable to conclude whether a specific user was in the input dataset. DP has emerged as the gold standard of data privacy. It has been extensively studied in various domains with the goal of safe-guarding sensitive data from adversarial attacks (Dwork et al.,, 2014; Abadi et al.,, 2016; McMahan et al.,, 2018). Impact of DP and designing efficient DP-preserving algorithms have also been studied for different settings of bandits, such as stochastic bandits (Mishra and Thakurta,, 2015; Tossou and Dimitrakakis,, 2016; Sajed and Sheffet,, 2019; Hu and Hegde,, 2022; Azize and Basu, 2022a, ), adversarial bandits (Tossou and Dimitrakakis,, 2017) and LCB (Shariff and Sheffet,, 2018; Neel and Roth,, 2018; Hanna et al.,, 2024; Azize and Basu,, 2024).
@@
  1. 1.
    What is the cost of ensuring $(\epsilon,\delta)$-JDP in high-dimensional sparse linear contextual bandits?
  2. 2.
    Can we design a privacy-preserving SLCB algorithm that achieves the minimal cost due to $(\epsilon,\delta)$-JDP?
@@
2. Algorithm and Regret Analysis: We propose a novel SLCB algorithm, namely FLIPHAT, and prove its $(\epsilon,\delta)$-JDP guarantee. We analyze the expected regret for FLIPHAT and demonstrate that in the problem-independent case, the upper bound matches our lower bound up to $s^{*}$ and logarithmic factors in $T$. We also numerically demonstrate its logarithmic dependence on $d$ across privacy regimes.
@@
4. Improved analysis of N-IHT: We provide a refined analysis for Peeling-based (Dwork et al.,, 2021) empirical risk minimization with or without a linear model. This yields a tighter bound on estimation error of N-IHT and is used in the regret analysis of FLIPHAT. The results and tools developed here might be of independent interest and can be applied in settings where N-IHT is deployed.
2 Problem Formulation: Sparse Linear Contextual Bandit with JDP
@@
Table 1: Comparison with prior works on private LCB (without margin condition)
@@
3 Lower Bound on Regret of SLCBs with JDP
@@
We define the minimax regret in the $(\epsilon,\delta)$-JDP SLCB as $R^{\mathrm{minimax}}_{(\epsilon,\delta)}(T):=\inf_{\pi\in\Pi(\epsilon,\delta)}\sup_{P_{\mathbf{X},\mathbf{r}}\in\mathcal{B}}\mathbb{E}[R_{\pi}(T)]$, where the supremum is taken over $\mathcal{B}$, the space of SLCB instances such that $\mathbb{E}\left\|x_{i}(t)\right\|_{\infty}\asymp 1$ for all arms $i$ and time $t$, and the infimum is taken over $\Pi(\epsilon,\delta)$, the space of $(\epsilon,\delta)$-JDP policies.
@@
If $\epsilon,\delta>0$ such that $\epsilon^{2}<\log(1/\delta)$, then for $d>9$ sufficiently large value for $s^{*}\log(d/s^{*})$ the minimax regret for SLCBs under $(\epsilon,\delta)$-JDP
$R^{\mathrm{minimax}}_{(\epsilon,\delta)}(T)=~{}\Omega\Big{(}\max\{\underbrace{\log(d/s^{*})\sqrt{s^{*}}\left\{\epsilon^{2}/\log(1/\delta)\right\}^{-1/2}}_{\text{private and high-d}},\underbrace{\sqrt{s^{*}T\log(d/s^{*})}}_{\text{non-private and high-d}}\}\Big{)}.$
Additionally, for $\delta=0$, we get
@@
Proof Technique. Our proof differs from the existing lower bound techniques for private bandits in (Shariff and Sheffet,, 2018; Basu et al., 2019b, ; Azize and Basu, 2022b, ; Azize and Basu,, 2024), and non-private SLCB setting (Li et al.,, 2021). Rather our proof depends on converting the regret lower bound problem to an estimation lower bound problem following He et al., (2022), and combing the information-theoretic techniques for sparse linear models (Duchi and Wainwright,, 2013) and privacy-restricted lower bound techniques in Acharya et al., (2021); Kamath et al., (2022).
4 Refined Analysis: Peeling for $(\epsilon,\delta)$-DP Empirical Risk Minimization
@@
The well-known Peeling algorithm (Algorithm 3) is used for $(\epsilon,\delta)$-DP top-$s$ selection problem (Dwork et al.,, 2021). We denote the output of Peeling applied to a vector $v\in\mathbb{R}^{d}$ as $\mathcal{P}_{s}(v;\epsilon,\delta,\lambda)$, where $\epsilon,\delta$ are privacy parameters and $\lambda$ is the scale parameter of the Laplace noise. For $v\in\mathbb{R}^{d}$, and an index set $S\subseteq[d]$, $v_{S}\in\mathbb{R}^{d}$ denotes the vector such that $v_{S}$ matches $v$ on indices in $S$ and is 0, elsewhere. Peeling returns a $s$-sparse vector in $\mathbb{R}^{d}$. It is important to point out that the peeling algorithm is an important building block of N-IHT, which in turn is a crucial component of FLIPHAT.
@@
Assume that $\left\|x_{i}\right\|_{\infty}\leq x_{\sf max}$ and $\theta^{*}\in\mathbb{R}^{d}$ is any point such that $\left\|\theta^{*}\right\|_{1}\leq b_{\sf max}$ and $\left\|\theta^{*}\right\|_{0}=s^{*}$. Now, we consider three good events. First, we consider the event
$\mathcal{E}_{1}:=\left\{\inf_{\theta}\phi_{\min}(Cs^{*},\nabla^{2}\mathcal{L}(\theta))\geq\underline{\kappa},\sup_{\theta}\phi_{\max}(Cs^{*},\nabla^{2}\mathcal{L}(\theta))\leq\bar{\kappa}\right\},$
where $\nabla^{2}\mathcal{L}$ is the Hessian computed for the data $(X,{\sf clip}_{R}(y))$. Second, we define $\mathcal{E}_{2}:=\left\{\left\|\nabla\mathcal{L}(\theta^{*})\right\|_{\infty}\leq\bar{g}\right\}$, where $\nabla\mathcal{L}$ is the corresponding gradient. Then finally, we define the event $\mathcal{E}_{3}:=\left\{\max_{m\in[M]}\mathbf{W}_{m}\leq\bar{W}_{(\epsilon,\delta)}\right\}.$ Here, $\mathbf{W}_{m}:=\sum_{i\in[s]}\|\mathbf{w}_{i}^{(m)}\|_{\infty}^{2}+\|\tilde{\mathbf{w}}_{S^{m}}^{(m)}\|_{2}^{2}$ is the total amount of noise injected in the $m$-th iteration (where $\mathbf{w}_{i}$ and $\tilde{\mathbf{w}}$ are the two sources of noise introduced by Peeling $\mathcal{P}_{s}$ in Line 3). With these events, we derive an upper bound on the $\ell_{2}$ distance of $\hat{\theta}$, i.e. the output of Algorithm 1, from $\theta^{*}$.
Algorithm 1 Noisy Iterative Hard Thresholding (N-IHT)
@@
Set $\eta=1/2\bar{\kappa},M\geq 28\kappa\log(b_{\sf max}^{2}n),C\geq\left\|\theta^{*}\right\|_{1}$ and $s$ $\asymp$ [^4] For two non-negative sequences $\{a_{n}\}_{n\in\mathbb{N}}$ and $\{b_{n}\}_{n\in\mathbb{N}}$, we use $a_{n}\asymp b_{n}$ to signify $a_{n}=O(b_{n})$ and $b_{n}=O(a_{n})$. $\kappa^{2}s^{*}$. Then, with probability at least $1-{\mathbb{P}}(\mathcal{E}_{1}^{c})-{\mathbb{P}}(\mathcal{E}_{2}^{c})-{\mathbb{P}}(\mathcal{E}_{3}^{c})$, we have
$\left\|\theta^{M}-\theta^{*}\right\|_{2}^{2}\leq\frac{1}{n}+C_{1}\kappa^{2}\bar{W}_{(\epsilon,\delta)}+C_{2}\frac{\kappa^{2}}{\underline{\kappa}^{2}}s^{*}\bar{g}^{2},$
for suitably large constants $C_{1},C_{2}>0$, where the probability measure ${\mathbb{P}}$ is that arising from the joint distribution $\mathcal{Q}$ and the randomness due to privacy. $\kappa=\bar{\kappa}/\underline{\kappa}$ denotes the condition number of the Hessian of the loss $\mathcal{L}$.
@@
5 FLIPHAT: Algorithm Design and Regret Analysis
@@ 5.1 Algorithm Design
To this end, we point out that FLIPHAT also enjoys the reward-DP (Hanna et al.,, 2024), where the algorithm only ensures the privacy protection of the reward stream. In this setup, the adversary is only allowed to perturb the reward stream which is quite a strong restriction on the nature of the adversary. JDP protects against a more general adversary and protects the information in both rewards and contexts. This is important for applications where context is a user data, e.g. personalized medicine.
@@ 5.2 Regret Analysis
In this section, we present the regret upper bound of FLIPHAT. We first briefly discuss the assumptions that we impose on the various model components, which are well accepted in the literature on SLCB (Zhang and Huang,, 2008; Zhang,, 2010; Bastani and Bayati,, 2020; Li et al.,, 2021; Chakraborty et al.,, 2023). In regret analysis, we do not require any additional assumption for privacy.
Definition 5.2 (Sparse Riesz Condition).
$\phi_{\min}(s;A):=\inf_{u:u\neq 0,\left\|u\right\|_{0}\leq s}u^{\top}Au/\left\|u\right\|_{2}^{2},\quad\phi_{\max}(s;A):=\sup_{u:u\neq 0,\left\|u\right\|_{0}\leq s}u^{\top}Au/\left\|u\right\|_{2}^{2}.$
@@
    Margin condition: There exists positive constants $\Delta_{*},A$ and $\alpha\in[0,\infty)$, such that for $h\in\left[A\sqrt{\log(d)/T},\Delta_{*}\right]$ and for all $t\in[T]$,
    ${\mathbb{P}}\left(x_{a^{*}_{t}}(t)^{\top}\beta^{*}\leq\max_{i\neq a^{*}_{t}}x_{i}(t)^{\top}\beta^{*}+h\right)\leq\left(\frac{h}{\Delta^{*}}\right)^{\alpha}.$
@@
Remarks: In Assumption 5.3(b), $\left\|X\right\|_{\psi_{2}}:=\sup_{u:\left\|u\right\|_{2}\leq 1}\left\|u^{\top}X\right\|_{\psi_{2}}$ is the Orlicz-norm, where for a random-variable $Z$, $\left\|Z\right\|_{\psi_{2}}:=\inf\{\lambda>0:\mathbb{E}\exp(Z^{2}/\lambda^{2})\leq 2\}$. It states that the context distribution for each arm is sub-Gaussian, this assumption does not require the distributions to have zero mean. We remark that while we assume i.i.d. contexts across time, for each time $t$, the contexts across different arms are allowed to be correlated. The zero-mean sub-Gaussian assumption on the noise (Assumption 5.5) is satisfied by various families of distributions, including normal distribution and bounded distributions, which are commonly chosen noise distributions. The margin condition (Assumption 5.4(b)) essentially controls the hardness of the bandit instance, with $\alpha=\infty$ making it easiest due to a deterministic gap between arms, and $\alpha=0$ making it hardest (no a-priori information about arm separation). A detailed discussion on the assumptions is in Appendix B.
Next, we derive a bound on estimation error of the parameter $\beta^{*}$ at any episode $\ell$, which is central to the regret analysis. Let us define the covariance matrix of the observed contexts of the $\ell$th episode as $\hat{\Sigma}_{\ell}=X_{\ell}X_{\ell}^{\top}/N_{\ell}$. The following result is a consequence of Theorem 4.1.
@@
Suppose Assumptions 5.3, 5.4(a) and 5.5 holds. If we set $\eta=1/2\bar{\kappa},C=b_{\sf max},s\asymp\kappa^{2}s^{*},M_{\ell+1}=28\kappa\log(b_{\sf max}^{2}N_{\ell})$ and $R=x_{\sf max}b_{\sf max}+\sigma\sqrt{2\log N_{\ell}}$, we have for all $\ell>1$
@@
with probability at least $1-c_{1}\exp(-c_{2}\log d)-{\mathbb{P}}(\mathcal{E}_{1\ell}^{c})-1/N_{\ell}$, where
$\vspace*{-1em}\Lambda_{\textsc{non-private}}$ $=\sigma^{2}\kappa^{2}\frac{s^{*}\log d}{\underline{\kappa}^{2}N_{\ell}},\quad\Lambda_{\textsc{private}}=\sigma^{2}\kappa^{2}\frac{(s^{*}\log d)^{2}\log\left(\frac{\log N_{\ell}}{\delta}\right)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}$
represent the non-private (statistical) and private parts of the estimation bound. In the above, $\mathcal{E}_{1\ell}=\{\underline{\kappa}\leq\phi_{\min}(Cs^{*},\hat{\Sigma}_{\ell})<\phi_{\max}(Cs^{*},\hat{\Sigma}_{\ell})\leq\bar{\kappa}\}$ and $\kappa=\bar{\kappa}/\underline{\kappa}$.
@@
Suppose Assumption 5.3, 5.4, and 5.5 hold. Then for any privacy level $\epsilon>0$ and $\delta<e^{-1}$, and with the same choices of the tuning parameters as in Proposition 5.6 with $\bar{\kappa}\asymp\log K$ and $\underline{\kappa}\asymp K^{-1}$, FLIPHAT enjoys the following regret bound
$\mathbb{E}(R(T))\lesssim x_{\sf max}b_{\sf max}{s^{*}}\log d+I_{\alpha}\vee J_{\alpha}(\epsilon,\delta),$
@@
$I_{\alpha}=\begin{cases}\frac{1}{\Delta_{*}^{\alpha}}\left(\frac{T^{\frac{1-\alpha}{2}-1}}{1-\alpha}\right)({s^{*}}^{2}\log d)^{\frac{1+\alpha}{2}},\\ {s^{*}}^{2}\log d,\\ \end{cases}J_{\alpha}(\epsilon,\delta)=\begin{cases}\Psi_{\alpha}(T)\left(\frac{{s^{*}}^{3}\log^{2}d\log(1/\delta)}{\epsilon^{2}}\right)^{\frac{1+\alpha}{2}},&\alpha\in[0,\infty)\\ \frac{{s^{*}}^{3}\log^{2}d\log(1/\delta)}{\epsilon^{2}},&\alpha=\infty.\end{cases}$

### v3

@@ Abstract
* Equal contribution.
@@ 1 INTRODUCTION
In many of these applications, the personal information used as contexts often has a large dimension, which can be modeled by viewing the contexts as high-dimensional vectors. This is relevant in the modern big data era, where the applications usually deal with many features. However, the established regret lower bound in LCB polynomially depends on the context dimension (Dani et al.,, 2008; Chu et al.,, 2011). Thus, additional structural conditions, such as sparsity, are needed to be enforced to make the problem tractable. Sparsity aligns with the idea that among these features, possibly only a small subset of them are relevant to generate the reward. This setting is called the high-dimensional sparse linear contextual bandit (SLCB). Recently, SLCB has received significant attention (Wang et al.,, 2018; Bastani and Bayati,, 2020; Hao et al.,, 2020; Chakraborty et al.,, 2023) and have been useful in practical applications (Miao and Chao,, 2022; Ren and Zhou,, 2024).
@@
Example 1. One of the earliest successful application of linear contextual bandits is personalized news article recommendation and its validation on Yahoo! Today Module by Li et al., (2010). In the dataset collected from Yahoo! platform, each user is represented by a 1193-dimensional features but for a successful deployment they are reduced to 5 dimensions. The feature of every user include their gender, age, geographical address, and even behavorial quantifiers. These are sensitive data as per GDPR (Voigt and Von dem Bussche,, 2017) and California Consumer Privacy Act (Harding et al.,, 2019). Thus, protecting data privacy in this problem, and designing efficient and private SLCB algorithms are fundamentally necessary.
Example 2. A recent example of deploying SLCB with user-data is optimizing click-through rates in the Tencent search advertising dataset (Tencent,, 2012). Though each context is 509256 dimensional, it lends itself to a sparse structure (Wang et al.,, 2023). Here, also the context incorporates sensitive information, like gender, age etc., which are protected by privacy laws. Thus, private SLCBs are a natural requirement here.
Example 3. Another application of SLCB is a mobile health application that recommends a personalized treatment plan (actions) to each patient (user) based on her personal information, such as age, weight, height, medical history etc. (Tewari and Murphy,, 2017; Killian et al.,, 2023). Here, for every user and personalized treatment plan contexts are generally obtained through a feature representation, where the dimension is often large but there are only fewer relevant features. The algorithm uses patients’ data including treatment outcomes and medical contexts of patients, which are sensitive information as per GDPR, CCPA, and HIPAA (Annas,, 2003). Hence, this is a natural application that motivates the study of efficient and privacy-preserving SLCB algorithms.
Related Work: Differential Privacy in Bandits. Problems like the ones in the aforementioned examples motivate us to study data-privacy preserving SLCB algorithms. Specifically, we follow Differential Privacy (DP) (Dwork,, 2006) as the data privacy framework. DP aims to ensure that while looking into the output of an algorithm, an adversary is unable to conclude whether a specific user was in the input dataset. DP has emerged as the gold standard of data privacy. It has been extensively studied in various domains with the goal of safe-guarding sensitive data from adversarial attacks (Dwork et al.,, 2014; Abadi et al.,, 2016; McMahan et al.,, 2018). Impact of DP and designing efficient DP-preserving algorithms have also been studied for different settings of bandits, such as stochastic bandits (Mishra and Thakurta,, 2015; Tossou and Dimitrakakis,, 2016; Sajed and Sheffet,, 2019; Hu and Hegde,, 2022; Azize and Basu, 2022a, ), adversarial bandits (Tossou and Dimitrakakis,, 2017) and LCB (Shariff and Sheffet,, 2018; Neel and Roth,, 2018; Hanna et al.,, 2024; Azize and Basu,, 2024).
@@
1. What is the cost of ensuring $(\epsilon,\delta)$-JDP in high-dimensional sparse linear contextual bandits?
2. Can we design a privacy-preserving SLCB algorithm that achieves the minimal cost due to $(\epsilon,\delta)$-JDP?
@@
2. Algorithm and Regret Analysis: We propose a novel SLCB algorithm, namely FLIPHAT, that uses N-IHT as a key component, and prove its $(\epsilon,\delta)$-JDP guarantee. We analyze the expected regret for FLIPHAT and demonstrate that in the problem-independent case, the upper bound matches our lower bound up to $s^{*}$ and logarithmic factors in $T$. We also numerically demonstrate its logarithmic dependence on $d$ across privacy regimes.
@@
4. Improved analysis of N-IHT: We provide a refined analysis for Peeling-based (Dwork et al.,, 2021) empirical risk minimization under non i.i.d. setting. This yields a bound on estimation error of N-IHT under the SLCB setting, and it is used in the regret analysis of FLIPHAT. The results and tools developed here might be of independent interest and can be applied in settings where N-IHT is deployed.
2 PROBLEM FORMULATION
@@
Table 1: Comparison with prior works on private LCB (under no-margin condition)
@@
3 LOWER BOUNDS ON REGRET
@@
We define the minimax regret in the $(\epsilon,\delta)$-JDP SLCB as $R^{\mathrm{minimax}}_{(\epsilon,\delta)}(T):=\inf_{\pi\in\Pi(\epsilon,\delta)}\sup_{P_{\mathbf{X},\mathbf{r}}\in\mathcal{B}}\mathbb{E}[R_{\pi}(T)]$, where the supremum is taken over $\mathcal{B}$, i.e. the space of SLCB instances, and the infimum is taken over $\Pi(\epsilon,\delta)$, i.e. the space of $(\epsilon,\delta)$-JDP policies.
@@
If $\epsilon,\delta>0$ and $\epsilon^{2}<\log(1/\delta)$, then for sufficiently large $s^{*}\log(d/s^{*})$ the minimax regret for SLCBs under $(\epsilon,\delta)$-JDP
$R^{\mathrm{minimax}}_{(\epsilon,\delta)}(T)=$
$\Omega\Big{(}\max\Big{\{}\underbrace{\sqrt{\frac{s^{*}\log^{2}(d/s^{*})\log(1/\delta)}{\epsilon^{2}}}}_{{\text{private}\text{and high-d}}},\underbrace{\sqrt{s^{*}T\log(d/s^{*})}}_{\begin{subarray}{c}\text{non-private}\\ \text{and high-d}\end{subarray}}\Big{\}}\Big{)}.$
Additionally, for $\delta=0$, we get $R^{\mathrm{minimax}}_{\epsilon}(T)=\Omega\Big{(}\max\{s^{*}\log^{3/2}(d/s^{*})\epsilon^{-1},{\sqrt{s^{*}T\log(d/s^{*})}}\}\Big{)}.$
@@
Proof Technique. Our proof differs from the existing lower bound techniques for private bandits in (Shariff and Sheffet,, 2018; Basu et al., 2019b, ; Azize and Basu, 2022b, ; Azize and Basu,, 2024), and non-private SLCB setting (Li et al.,, 2021). Rather our proof depends on converting the regret lower bound problem to an estimation lower bound problem following He et al., (2022), and combing the information-theoretic techniques for sparse linear models (Duchi and Wainwright,, 2013) and privacy-restricted lower bound techniques in (Azize and Basu,, 2024; Kamath et al.,, 2022).
4 REFINED ANALYSIS: PEELING FOR $(\epsilon,\delta)$-DP ERM
@@
The well-known Peeling algorithm (Appendix A) is used for $(\epsilon,\delta)$-DP top-$s$ selection problem (Dwork et al.,, 2021). We denote the output of Peeling applied to a vector $v\in\mathbb{R}^{d}$ as $\mathcal{P}_{s}(v;\epsilon,\delta,\lambda)$, where $\epsilon,\delta$ are privacy parameters and $\lambda$ is the scale parameter of the Laplace noise. For $v\in\mathbb{R}^{d}$, and an index set $S\subseteq[d]$, $v_{S}\in\mathbb{R}^{d}$ denotes the vector such that $v_{S}$ matches $v$ on indices in $S$ and is 0, elsewhere. Peeling returns a $s$-sparse vector in $\mathbb{R}^{d}$. It is important to point out that the peeling algorithm is an important building block of N-IHT, which in turn is a crucial component of FLIPHAT.
@@
Assume that $\left\|x_{i}\right\|_{\infty}\leq x_{\sf max}$ and $\theta^{*}\in\mathbb{R}^{d}$ is any point such that $\left\|\theta^{*}\right\|_{1}\leq b_{\sf max}$ and $\left\|\theta^{*}\right\|_{0}=s^{*}$. Now, we consider three good events. First, we define the event $\mathcal{E}_{1}$ as the one satisfying the infimum of minimum sparse eigenvalue, i.e. $\inf_{\theta}\phi_{\min}(Cs^{*},\nabla^{2}\mathcal{L}(\theta))\geq\underline{\kappa}$, and the supremum of maximum sparse eigenvalue, i.e. $\sup_{\theta}\phi_{\max}(Cs^{*},\nabla^{2}\mathcal{L}(\theta))\leq\bar{\kappa}$ (ref. Definition B.1). Here, $\nabla^{2}\mathcal{L}$ is the Hessian computed for the data $(X,{\sf clip}_{R}(y))$. Second, $\mathcal{E}_{2}:=\left\{\left\|\nabla\mathcal{L}(\theta^{*})\right\|_{\infty}\leq\bar{g}\right\}$, where $\nabla\mathcal{L}$ is the corresponding gradient. Finally, $\mathcal{E}_{3}:=\left\{\max_{m\in[M]}\mathbf{W}_{m}\leq\bar{W}_{(\epsilon,\delta)}\right\}.$ Here, $\mathbf{W}_{m}:=\sum_{i\in[s]}\|\mathbf{w}_{i}^{(m)}\|_{\infty}^{2}+\|\tilde{\mathbf{w}}_{S^{m}}^{(m)}\|_{2}^{2}$ is the total amount of noise injected in the $m$-th iteration, where $\mathbf{w}_{i}$ and $\tilde{\mathbf{w}}$ are the two sources of noise introduced by Peeling $\mathcal{P}_{s}$ in Algorithm 3. With these events, we derive an upper bound on the $\ell_{2}$ distance of $\hat{\theta}$, i.e. the output of Algorithm 1, from $\theta^{*}$.
Algorithm 1 N-IHT
@@
Let $\eta=1/2\bar{\kappa},M\geq 28\kappa\log(b_{\sf max}^{2}n)$, $C\geq\left\|\theta^{*}\right\|_{1}$ and $s\gtrsim\kappa^{2}s^{*}$. Then, with probability at least $1-{\mathbb{P}}(\mathcal{E}_{1}^{c})-{\mathbb{P}}(\mathcal{E}_{2}^{c})-{\mathbb{P}}(\mathcal{E}_{3}^{c})$,
$\left\|\theta^{M}-\theta^{*}\right\|_{2}^{2}\leq\frac{1}{n}+C_{1}\kappa^{2}\bar{W}_{(\epsilon,\delta)}+C_{2}\frac{\kappa^{2}}{\underline{\kappa}^{2}}(s+s^{*})\bar{g}^{2},$
for suitably large constants $C_{1},C_{2}>0$, where the probability measure ${\mathbb{P}}$ is that arising from the joint distribution $\mathcal{Q}$ and the randomness due to privacy. $\kappa=\bar{\kappa}/\underline{\kappa}$ denotes the condition number of the Hessian of loss $\mathcal{L}$.
@@
5 FLIPHAT: ALGORITHM DESIGN & REGRET ANALYSIS
@@ 5.1 Algorithm Design
To this end, we point out that FLIPHAT also enjoys the reward-DP (Hanna et al.,, 2024), where the algorithm only ensures the privacy protection of the reward stream. In this setup, the adversary is only allowed to perturb the reward stream which is quite a strong restriction on the nature of the adversary. JDP protects against a more general adversary and protects the information in both rewards and contexts.
@@ 5.2 Regret Analysis
Now, we present the regret upper bound of FLIPHAT. We first briefly discuss the assumptions that we impose on the various model components, which are well accepted in the literature on SLCB (Zhang and Huang,, 2008; Zhang,, 2010; Bastani and Bayati,, 2020; Li et al.,, 2021; Chakraborty et al.,, 2023). In regret analysis, we do not require any additional assumption for privacy.
@@
    Margin condition: There exists positive constants $\Delta_{*},A$ and $\alpha\in[0,\infty)$, such that for $h\in\left[A\sqrt{\log(d)/T},\Delta_{*}\right]$ and $\forall t\in[T]$, ${\mathbb{P}}\left(x_{a^{*}_{t}}(t)^{\top}\beta^{*}\leq\max_{i\neq a^{*}_{t}}x_{i}(t)^{\top}\beta^{*}+h\right)\leq\left(\frac{h}{\Delta^{*}}\right)^{\alpha}$
@@
Remarks: In  5.2(b), $\left\|X\right\|_{\psi_{2}}:=\sup_{u:\left\|u\right\|_{2}\leq 1}\left\|u^{\top}X\right\|_{\psi_{2}}$ is the Orlicz-norm, where for a random-variable $Z$, $\left\|Z\right\|_{\psi_{2}}:=\inf\{\lambda>0:\mathbb{E}\exp(Z^{2}/\lambda^{2})\leq 2\}$. It states that the context distribution for each arm is sub-Gaussian, this assumption does not require the distributions to have zero mean. We remark that while we assume i.i.d. contexts across time, for each time $t$, the contexts across different arms are allowed to be correlated. The zero-mean sub-Gaussian assumption on the noise (Assumption 5.4) is satisfied by various families of distributions, including normal distribution and bounded distributions, which are commonly chosen noise distributions. The margin condition (Assumption 5.3(b)) essentially controls the hardness of the bandit instance, with $\alpha=\infty$ making it easiest due to a deterministic gap between arms, and $\alpha=0$ making it hardest, i.e. no apriori information about arm separation. A detailed discussion on the assumptions is in Appendix B.
Next, we derive a bound on estimation error of the parameter $\beta^{*}$ at any episode $\ell$, which is central to the regret analysis and a consequence of Theorem 4.1. Note that $(X_{\ell},y_{\ell})$ are only conditionally i.i.d., and not marginally independent.
@@
Suppose Assumptions 5.2, 5.3(a) and 5.4 holds. If we set $\eta=1/2\bar{\kappa},C=b_{\sf max},s\gtrsim\kappa^{2}s^{*},M_{\ell+1}=28\kappa\log(b_{\sf max}^{2}N_{\ell})$ and $R=x_{\sf max}b_{\sf max}+\sigma\sqrt{2\log N_{\ell}}$, we have for all $\ell>1$
@@
with probability at least $1-c_{1}\exp(-c_{2}\log d)-{\mathbb{P}}(\mathcal{E}_{1\ell}^{c})-1/N_{\ell}$, where $\Lambda_{\textsc{non-private}}=\sigma^{2}\kappa^{2}\frac{(s+s^{*})\log d}{\underline{\kappa}^{2}N_{\ell}}$, and $\Lambda_{\textsc{private}}=\sigma^{2}\kappa^{2}\frac{(s\log d)^{2}\log\left(\frac{\log N_{\ell}}{\delta}\right)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}$ are the non-private (statistical) and private terms in the estimation bound. Here, $\mathcal{E}_{1\ell}=\{\underline{\kappa}\leq\phi_{\min}(Cs^{*},\hat{\Sigma}_{\ell})<\phi_{\max}(Cs^{*},\hat{\Sigma}_{\ell})\leq\bar{\kappa}\}$, $\kappa=\bar{\kappa}/\underline{\kappa}$, and $\hat{\Sigma}_{\ell}=X_{\ell}X_{\ell}^{\top}/N_{\ell}$ is the covariance of the contexts observed in $\ell$-th episode.
@@
Suppose Assumption 5.2, 5.3, and 5.4 hold. Then for any privacy level $\epsilon>0$ and $\delta<e^{-1}$, and with the same choices of the tuning parameters as in Proposition 5.5 with $\bar{\kappa}\asymp\log K$ and $\underline{\kappa}\asymp K^{-1}$, FLIPHAT enjoys the following regret bound
$\mathbb{E}[R(T)]\lesssim x_{\sf max}b_{\sf max}(s+s^{*})\log d+I_{\alpha}\vee J_{\alpha}(\epsilon,\delta),$
@@
$I_{\alpha}=\begin{cases}\frac{1}{\Delta_{*}^{\alpha}}\left(\frac{T^{\frac{1-\alpha}{2}-1}}{1-\alpha}\right)\{(s+s^{*})^{2}\log d\}^{\frac{1+\alpha}{2}},&\alpha\in[0,\infty)\\ (s+s^{*})^{2}\log d,&\alpha=\infty.\\ \end{cases}$
$J_{\alpha}(\epsilon,\delta)=\begin{cases}\Psi_{\alpha}(T)\left\{\frac{(s+s^{*})^{3}\log^{2}d\log(1/\delta)}{\epsilon^{2}}\right\}^{\frac{1+\alpha}{2}},&\alpha\in[0,\infty)\\ \frac{(s+s^{*})^{3}\log^{2}d\log(1/\delta)}{\epsilon^{2}},&\alpha=\infty.\end{cases}$